OpenStudy (anonymous):

solve by elimination way 2x-y=8 and 2x+y=6

OpenStudy (anonymous):

add both equations and you get 4x = 14 x = 3.5 and plug this into first equation to get 7 - y = 8 y = -1 answer (3.5, -1)

OpenStudy (anonymous):

ok daverer?

OpenStudy (anonymous):

or have u any questions

OpenStudy (anonymous):

how did you do these i am new to these.....

OpenStudy (anonymous):

by the way sorry for the late reply

OpenStudy (anonymous):

ok - but i have to eat now - be back in 10 minutes

OpenStudy (anonymous):

okay.no problem i'll wait

OpenStudy (anonymous):

sorry i was longer than i thoght id be 2x-y=8 2x+y=6 there are two unknowns but can be solved as we have 2 equations we can add or subtract equations if we add the above the y disappears because -y + y = 0 so we get 4x + 0 = 14 4x = 14 divide both sides by 4 : x = 14/4 = 3.5 Now we can plug x = 3.5 into the second equation (any one will do): 2(3.5) + y = 6 7 + y = 6 y = 6-7 = -1 so thats it x=3.5 and y =-1 this is the solution we can now check if these are correct by substituting our vales for x and y into the first equation 2x - y = 8 so 2(3.5) - (-1) = 7 + 1 = 8 whch checls out